# Unveiling the Mathematical Reasoning in DeepSeek Models: A Comparative Study of Large Language Models

### v1

@@ 3.2 Datasets
To evaluate the mathematical reasoning capabilities of LLMs, we employ well-established benchmark datasets that encompass diverse mathematical domains, question types, and difficulty levels. The selected datasets include Math Competition (MATH) [40], Grade School Math 8K (GSM8K)  [16], and Massive Multitask Language Understanding (MMLU) [39], covering topics ranging from elementary arithmetic to advanced algebra, logic, and mathematical competitions. From each benchmark, we carefully select a representative subset of problems to ensure a balanced evaluation while maintaining the diversity and complexity of mathematical reasoning tasks. These subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ symbolic manipulation, logical reasoning, and problem-solving abilities [35, 74, 101]. In total, our evaluation spans 2,178 mathematical problems across multiple benchmark datasets, ensuring a comprehensive assessment of LLMs’ mathematical reasoning capabilities.
By evaluating LLMs on these benchmarks, we identified their strengths and limitations in mathematical reasoning, aiding in the development of more robust and capable models. Table 1 provides an overview of the datasets utilized in this study.
@@
3.2.1 MATH
MATH is a comprehensive benchmark dataset designed to evaluate mathematical reasoning across a wide range of problem-solving scenarios [40]. Unlike conventional mathematics problem datasets, which primarily assess algorithmic proficiency, MATH focuses on advanced reasoning and heuristic-driven problem-solving.
Specifically, MATH is derived from high-level mathematics competitions, including the AMC 10, AMC 12, and AIME [40]. These problems are inherently more complex than standard K-12 mathematics tasks, requiring the application of non-trivial problem-solving strategies, logical inference, and domain-specific heuristics rather than direct formulaic computations. Meanwhile, MATH is widely regarded as a difficult benchmark, with LLMs achieving accuracy rates between 3.0% and 6.9% [40]. Despite these low overall scores, LLMs demonstrate some degree of mathematical competency, as evidenced by: Up to 15% accuracy on the easiest difficulty level [40]. The ability to generate step-by-step solutions that, while sometimes incorrect, remain coherent and contextually relevant. To gauge the dataset’s difficulty, human performance was also considered. For instance, a computer science Ph.D. student with no particular affinity for mathematics achieved approximately 40% accuracy. In addition, a three-time IMO (International Mathematical Olympiad) gold medalist attained 90% accuracy. Thus, these findings indicate that MATH is challenging for both human solvers and LLMs, making it an invaluable dataset for assessing and advancing mathematical reasoning capabilities in both LLMs and human problem solvers.
3.2.2 GSM8K
GSM8K, introduced in 2021, is a benchmark dataset consisting of 8,500 high-quality grade school-level math problems  [16]. Designed to incorporate high linguistic diversity while relying on fundamental mathematical concepts, it presents a unique challenge for state-of-the-art LLMs. Although the underlying math is relatively simple, the diverse problem formulations create significant hurdles, preventing many models from achieving consistently high accuracy. GSM8K is structured into 7,500 training problems and 1,320 testing problems, all carefully crafted by expert human problem writers. The problems primarily involve elementary arithmetic operations and typically require 2 to 8 logical steps to reach a solution. This dataset is widely used to evaluate logical reasoning and mathematical proficiency in LLMs and serves as a benchmark for various assessments, including the LLM Leaderboard. Notably, while some GSM8K problems are conceptually straightforward, they can still be challenging for even the most advanced LLMs, often exhibiting high variability in responses when the same problem is presented in slightly different ways. This highlights the ongoing difficulty of achieving robust mathematical reasoning in language models. In this study, we selected the 1.32K mathematical test set from GSM8K to systematically evaluate and compare the mathematical reasoning capabilities of each LLM.
3.2.3 MMLU
The MMLU dataset is a comprehensive benchmark comprising multiple-choice questions from a diverse range of academic disciplines. Spanning 57 distinct tasks, it covers subjects in the humanities, social sciences, hard sciences, and mathematical reasoning, reflecting the breadth of knowledge that is essential for various fields of study [39]. MMLU’s questions were manually curated by graduate and undergraduate students from publicly available sources, including practice questions from standardized exams such as the Graduate Record Examination (GRE) and the United States Medical Licensing Examination (USMLE). Additionally, it features questions designed for undergraduate courses and readers of Oxford University Press books. Notably, its mathematical reasoning component encompasses multiple subfields, categorized into areas such as "Abstract Algebra," "College Mathematics," "Formal Logic," and "High School Mathematics." For instance, questions in the "Abstract Algebra" category originate from professional mathematical practice, while the "High School Mathematics" category includes problems akin to those found in standard high school exams. In total, MMLU comprises 15,908 questions, divided into three subsets: a). A few-shot development set containing five questions per subject. b). A validation set with 1,540 questions, used for selecting hyperparameters. c). A test set with 14,079 questions, ensuring a rigorous evaluation Each subject within MMLU includes a minimum of 100 test questions, making it a more extensive and challenging assessment than most standard exams. This broad and structured dataset serves as a critical benchmark for evaluating the reasoning and problem-solving capabilities of LLMs.
3.3 Evaluate LLMs in Mathematical Reasoning
As discussed previously, multiple LLMs have demonstrated remarkable abilities in mathematical reasoning [105]. To systematically evaluate the mathematical reasoning capabilities of various LLMs, we conducted experiments using API-based access to a diverse set of models. The selected models include Google’s Gemini-2.0-flash, OpenAI’s GPT-4o, o1-mini, o1, and o3-mini, and DeepSeek’s DeepSeek-1.5B and DeepSeek-R1. These models were accessed through their respective API services using secure API keys from Google, OpenAI, DeepSeek, and Openrouter.
3.3.1 Prompting Strategy
The effectiveness of LLMs in mathematical reasoning is highly dependent on the design of input prompts [5, 10, 25]. To ensure a fair and reproducible evaluation, we employed a structured prompting strategy focusing on zero-shot prompting initially. Our methodology was designed to assess models’ inherent problem-solving abilities without additional external guidance [87]. Overall, for all evaluations, we incorporated a zero-shot prompting approach, where models were presented with mathematical problems without any prior exemplars or demonstrations. Zero-shot prompting is particularly valuable in assessing a model’s ability to generalize mathematical knowledge from its pretraining corpus and apply learned concepts to novel problems [87, 98]. By not providing explicit examples, this method evaluates how well a model can infer the appropriate solution methodology based solely on its prior learning.

### v2

@@ 2.2 Datasets
To evaluate the mathematical reasoning capabilities of LLMs, we employ well-established benchmark datasets that encompass diverse mathematical domains, question types, and difficulty levels. The selected datasets include Math Competition (MATH) [Hendrycks et al., 2021], Grade School Math (GSM8K) [Cobbe et al., 2021], and Massive Multitask Language Understanding (MMLU) [Hendrycks et al., 2020] math subset, covering topics ranging from elementary arithmetic to advanced algebra, logic, and mathematical competitions. From each benchmark, we carefully select a representative subset of problems to ensure a balanced evaluation while maintaining the diversity and complexity of mathematical reasoning tasks. These subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ symbolic manipulation, logical reasoning, and problem-solving abilities [Gema et al., 2024, Setlur et al., 2024, Zhang et al., 2024].
Figure 2: Data distribution across selected datasets for evaluating mathematical reasoning.
Notably, these subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ logical reasoning, and problem-solving capabilities [Gema et al., 2024, Setlur et al., 2024, Zhang et al., 2024]. We selected 262 free-form numeric problems from MATH dataset (mathematics competitions), covering domains such as algebra, geometry, number theory, and combinatorics. For this study, we selected problems spanning $\geq 3$ difficulty levels to ensure coverage between medium and hard tasks. The GSM8K testset includes 1,320 arithmetic word problems designed for the grade school level. From MMLU, we incorporate four subject-specific subsets—College Mathematics (MCM), Abstract Algebra (MAA), Formal Logic (MFL), and High School Mathematics (MHM)—ranging from 100 to 270 multiple-choice questions. These subsets cover diverse topics such as calculus, group theory, logical reasoning, and general high school math. In total, our experimental benchmark includes 2,178 problems, offering broad coverage across difficulty levels, formats, and mathematical domains to support fine-grained evaluation of reasoning skills in LLMs. By evaluating LLMs on these benchmarks, we identified their strengths and limitations in mathematical reasoning, aiding in the development of more robust and capable models Figure 2 presents an overview structure of the datasets utilized in this study. The detailed table of the dataset is shown in Appendix A.
2.3 Evaluation of Mathematical Reasoning Across LLMs
As discussed previously, multiple LLMs have demonstrated remarkable capabilities in mathematical reasoning [Zhou et al., 2024]. To systematically and comprehensively evaluate the mathematical reasoning capabilities of various LLMs, we conducted experiments using API-based access to a diverse set of models. The selected models include OpenAI’s GPT-4o, o1-mini, o1, and o3-mini, DeepSeek’s DeepSeek-1.5B and DeepSeek-R1, Google’s Gemini 2.0 Flash, and Meta’s Llama-4 the latest one. These models were accessed through their respective API services using secure API keys from Google, OpenAI, DeepSeek, and Openrouter.
2.3.1 Prompting Strategy
The effectiveness of LLMs in mathematical reasoning is highly dependent on the design of input prompts [Amatriain, 2024, Cain, 2024, Denny et al., 2024]. To ensure a fair and reproducible evaluation, we employed a structured prompting strategy focusing on zero-shot prompting initially. Our methodology was designed to assess models’ inherent problem-solving abilities without additional external guidance [Wu et al., 2024]. Overall, for all evaluations, we incorporated a zero-shot prompting approach, where models were presented with mathematical problems without any prior exemplars or demonstrations. Zero-shot prompting is particularly valuable in assessing a model’s ability to generalize mathematical knowledge from its pretraining corpus and apply learned concepts to novel problems [Wu et al., 2024, Yuan et al., 2024]. By not providing explicit examples, this method evaluates how well a model can infer the appropriate solution methodology based solely on its prior learning.
For multiple-choice questions, we provided the set of answer choices and instructed the models to select the correct response from the given options, ensuring a structured decision-making process [Myrzakhan et al., 2024]. In contrast, for open-ended problems that required free-form numeric or symbolic responses, we presented the question directly without predefined answer choices, allowing the models to generate responses based on their internal reasoning capabilities [Tam et al., 2025]. This approach ensured that multiple-choice tasks evaluated the models’ ability to differentiate between structured options. The prompt templates linked with each category question are elaborated in Appendix B.
2.3.2 Settings
To ensure a standardized evaluation across different LLMs, we utilized API access through OpenAI and OpenRouter, while maintaining their respective default generation settings. All models, including GPT-4o and o-series models, Gemini, Llama-4 and DeepSeek use a default temperature of $T=1.0$ and an unrestricted probability sampling strategy (top-p = 1.0) [Nguyen et al., 2024], allowing the model to consider the full probability distribution for token selection.
2.3.3 Evaluation Metric
We evaluate the model outputs using three core metrics, Correctness, Clarity, and Reasoning, which reflect different dimensions of mathematical understanding and communication. The evaluation is conducted using LLM-as-a-judge methodology implemented via the DeepEval framework.[Ip and Vongthongsri, 2025]
Correctness measures whether the model output reaches the right final answer or conveys the correct mathematical content, allowing for variations in formatting. Clarity assesses the logical coherence and readability of the explanation, ensuring that the reasoning process is easy to follow. Reasoning evaluates the depth and validity of the model’s step-by-step problem-solving approach.
Each model-generated response is processed through a pre-defined prompt template with access to the original question and reference answer. DeepEval internally prompts a trusted evaluator model (e.g., GPT-4o) to return scalar ratings for each metric. Let $M$ be a model, $Q$ be a set of questions, and $\text{Score}_{M}^{(k)}(q)$ be the score for metric $k\in\{\text{Correctness},\text{Clarity},\text{Reasoning}\}$ on question $q\in Q$. The average metric score is computed as:
$\text{Score}_{M}^{(k)}=\frac{1}{|Q|}\sum_{q\in Q}\text{LLM}_{\text{eval}}\left(\text{prompt}(q,\text{output}_{M}(q),\text{gold}(q),\text{steps}_{k})\right)$
3 Results and Analysis
In this section, we conducted a quantitative and qualitative evaluation of the state-of-the-art LLMs mentioned earlier for mathematical reasoning. As an initial assessment, we selected a subset of three widely recognized datasets [Cobbe et al., 2021, Hendrycks et al., 2020, Lu et al., 2023], as introduced in Section 3, to systematically evaluate the performance of various cutting-edge models. Overall, our evaluation in Table 3, covering diverse evaluation metric, such as Accuracy(Acc), Reasoning Score(RS), and Clarity Score(CS), includes DeepSeek-R1 [Liu et al., 2024a],along with its distilled variant DeepSeek-1.5B [Guo et al., 2025], which is derived from the R1 model. Additionally, we assessed Google’s latest Gemini 2.0 Flash [Imran and Almusharraf, 2024], four OpenAI models: GPT-4o, o1-mini, o1, and o3-mini [Arrieta et al., 2025, Hurst et al., 2024, Zhong et al., 2024] and finally Meta’s Llama-4 the latest model [AI, 2025]. Notably, while all these models are designed with reasoning capabilities, GPT-4o and Llama-4 are not explicitly optimized for providing reasoning-based CoT responses but they can still perform CoT reasoning quite well when prompted, especially in instruction-tuned variants. These selected LLMs represent the most advanced AI systems available today, continuously evolving and competing across various mathematical reasoning benchmarks.
Table 1: Performance comparison of LLMs across multiple mathematical reasoning abilities using three metrics: Accuracy (ACC), Reasoning Score (RS), and Clarity Score (CS). The results highlight the average performance of each model across all categories.
LLMs Math GSM8K MAA MFL MCM MHM Avg.
[…]
Deepseek-R1 92.37 96.36 95.00 97.62 96.00 97.78 95.75
Deepseek-1.5B 68.70 81.58 61.00 52.38 66.00 88.15 67.25
o1 95.80 96.21 94.00 96.51 99.00 99.26 96.91
o1-mini 89.69 96.13 88.00 78.57 89.29 98.89 88.89
o3-mini 96.56 94.84 96.00 95.24 99.00 98.52 97.06
Gemini-2.0-Flash 87.02 94.24 88.00 90.48 96.00 97.77 91.85
Llama-4 81.68 97.04 88.00 90.48 93.00 97.41 90.11
[…]
Deepseek-R1 95.15 98.15 94.76 94.04 98.26 98.19 96.08
Deepseek-1.5B 77.58 88.56 72.54 66.95 80.82 89.33 77.64
o1 98.73 98.82 94.00 86.84 99.09 98.20 95.97
o1-mini 95.51 98.46 87.49 95.63 93.09 98.51 94.04
o3-mini 99.04 97.53 97.37 98.44 99.05 98.33 98.45
Gemini-2.0-Flash 89.52 96.01 93.54 93.78 92.07 97.12 93.60
Llama-4 92.34 98.15 93.52 95.59 95.54 97.99 94.59
[…]
Deepseek-R1 85.89 97.21 87.81 89.13 84.42 88.10 87.07
Deepseek-1.5B 77.21 88.33 60.09 66.27 80.27 83.27 73.62
o1 86.80 98.34 84.91 77.76 85.00 87.95 84.48
o1-mini 87.87 98.11 80.52 89.64 81.29 88.25 85.11
o3-mini 87.43 96.88 89.89 91.55 91.26 88.26 89.07
Gemini-2.0-Flash 80.41 90.48 81.93 88.13 79.98 85.23 83.14
Llama-4 87.73 94.71 81.79 89.85 85.69 89.07 86.43
3.1 Overview of Benchmark Performance
We begin our evaluation with a comprehensive comparison of overall accuracy across the benchmark tasks, including MATH, GSM8K, and four mathematical subsets from the MMLU corpus(MAA, MFL, MCM, MHM). Table 3 summarizes the accuracy score (ACC), clarity score (CS)and reasoning scores (RS), highlighting each model’s performance and robustness across these domains.
Overall, the o3-mini model achieved the highest average accuracy, outperforming all baselines with an average of 97.06%. It showed particularly strong performance in the Math (96.56%), MAA (96.00%), and MCM (99.00%) subsets. Similarly, DeepSeek-R1 and o1 demonstrated competitive accuracy, achieving 95.75% and 96.91% respectively. When evaluating reasoning score, o3-mini again led the comparison with an average score of 98.45%, closely followed by DeepSeek-R1 (96.08%) and o1 (95.97%). These models consistently showed high reasoning coherence and correctness across all mathematical domains. Importantly, we include representative examples of both correct and incorrect responses in Appendix C.
In terms of clarity score, which evaluates the explanation’s readability and logical structure, the results followed a similar trend. o1 and o1-mini scored highest with averages of 88.44% and 85.11%, respectively. Notably, GPT-4o, although not optimized explicitly for math reasoning, maintained competitive scores in clarity (87.33%) and reasoning (94.32%).
Notably, the distilled variant DeepSeek-1.5B exhibited markedly lower scores across all three metrics, especially in reasoning (77.64%) and clarity (73.62%), emphasizing the performance compromises introduced by model compression. Appendix D provides a detailed comparative breakdown of each model’s performance.
In summary, while all evaluated models show proficiency, our findings reveal significant variance in cross-domain generalization. The superior performance of models like DeepSeek-R1 and o1 across diverse mathematical tasks positions them as current leaders in LLM-based mathematical reasoning, while others specially deepseek’s distilled variant, despite it’s scale, remain constrained by limitations in symbolic abstraction and formal manipulation.
3.2 Cognition Profile Across Reasoning Dimensions
To better understand the model-wise variation in mathematical reasoning behavior, we visualize the cognition profiles of each evaluated LLM using radar plots (Figure 3). Each subplot corresponds to a single model, showing its relative performance across six domains: MATH, GSM8K, MAA, MFL, MCM, and MHM. The colored lines represent the evaluation dimensions—Accuracy, Reasoning, and Clarity—as measured by the DeepEval framework [Ip and Vongthongsri, 2025]. The Figure 3 demonstrates, o3-mini demonstrates consistent and high performance across all dimensions, with particularly dominant accuracy and reasoning scores, forming a near-regular and saturated polygon. This consistency highlights the effectiveness of its architecture in capturing logical structure and factual correctness.
Figure 3: Radar plots of LLM performance across six mathematical domains. Each sub-plot corresponds to a model. Lines represent three evaluation metrics: Accuracy, Reasoning, and Clarity. This visualization highlights the strengths and weaknesses of each model across different reasoning dimensions and content types.
In contrast, the DeepSeek-1.5B (distilled) model displays notable asymmetry in its radar shape, indicating uneven performance across both domains and evaluation metrics. Specifically, the reduced reasoning and clarity scores in the MAA and MFL domains suggest limitations in handling abstract and symbolic reasoning, a likely result of aggressive distillation. On the other hand, GPT-4o model shows strong reasoning capability, especially in GSM8K and MFL, though with relatively lower accuracy in the MATH domain. Its clarity curve remains stable, affirming its general-purpose alignment rather than domain-specific tuning. DeepSeek-R1 and o1 both display high and balanced performance, especially in the reasoning dimension, validating their suitability for high-precision mathematical applications. Meanwhile, Llama-4 and Gemini 2.0 Flash show strong clarity but slightly lower reasoning compared to the top-tier models.
Overall, these cognition profiles emphasize that high performance in mathematical reasoning tasks depends not just on accuracy but also on the model’s ability to maintain logical flow and human-readable clarity across domains. The radar plots provide an interpretable and multi-dimensional view into how different models prioritize or trade off among these cognitive capabilities.
3.3 Accuracy–Latency Trade-off Analysis
Figure 4: Mean accuracy (%) versus average response time (seconds) for each model. The bar height indicates average accuracy across reasoning tasks, while the color gradient and labels represent the average response latency.
To complement our evaluation of cognitive performance, we investigate the trade-off between reasoning accuracy and computational efficiency across models (Figure 4). Each bar represents a model’s average accuracy across our mathematical reasoning benchmarks, while the color gradient and numerical labels denote the average response time (in seconds), measured using consistent inference settings.
Among the models, o3-mini achieves an optimal balance, with a high accuracy of 97.1% and a relatively low response time of 6.8 seconds. This performance places it at the efficiency frontier, demonstrating both computational effectiveness and strong reasoning capabilities. Similarly, o1-mini maintains respectable accuracy (88.9%) with an even lower latency of 5.3 seconds, making it highly suitable for real-time or resource-constrained deployments.
Furthermore, Gemini 2.0 Flash stands out for its minimal latency of just 4.2 seconds while achieving a solid 91.8% accuracy. This profile makes it particularly attractive for latency-sensitive applications. In contrast, models like o1 and DeepSeek-R1 deliver exceptional accuracy scores (96.9% and 95.8%, respectively) but exhibit significantly longer response times (15.0 and 81.05 seconds), which may present challenges in time-critical settings. The DeepSeek-1.5B model highlights the trade-off at the other end of the spectrum, with fast inference (5.0 seconds) but substantially lower accuracy, reflecting the compromises often seen in smaller or distilled architectures. Overall, this comparison underscores a clear trade-off between accuracy and inference speed. While some models prioritize correctness at the cost of latency, others offer rapid responses with modest performance degradation.
4 Discussion
In this study, we conducted a comprehensive evaluation of various LLMs to assess their mathematical reasoning capabilities across multiple benchmarks [Cobbe et al., 2021, Hendrycks et al., 2020, 2021].Our empirical analysis reveals that o3-mini outperformed all evaluated models in terms of overall accuracy and reasoning quality across a range of mathematical domains. It demonstrated particular strength in complex and abstract areas, such as Abstract Algebra and Multistep Calculation, while maintaining relatively low latency. In comparison, DeepSeek-R1 and o1 also exhibited strong reasoning capabilities, especially in structured arithmetic domains, but their performance was slightly lower than o3-mini in both accuracy and efficiency. These results confirm the existence of a trade-off between computational efficiency and reasoning performance, which must be carefully considered when designing LLMs for downstream applications.
Furthermore, our results validate the advanced design of DeepSeek-R1, which employs GRPO—a reinforcement learning-based optimization approach—instead of conventional supervised finetuning [Evstafev, 2025, Liu et al., 2024a, Neha and Bhati, 2025]. GRPO allows the model to evaluate its outputs relative to a group of sampled peers, eliminating the need for external reward models [Liu et al., 2024a, b]. This strategy achieves strong performance, particularly on GSM8K and logic-based tasks, it remains slightly behind o3-mini in accuracy and reasoning consistency across all domains. These comparisons suggest that reinforcement learning with GRPO enhances model generalization, but domain-specific prompting strategies—like those used in o3-mini—can still offer an edge in specialized reasoning tasks.
Moreover, proof-based reasoning, a crucial component of higher-level mathematics, demands meta-cognitive processes that extend beyond pattern recognition, including recursive logic application and theorem proving. The limitations observed in DeepSeek-R1 to generalize across abstract mathematical problems indicate that a more versatile reasoning approach may be needed to complement GRPO-based training. In contrast, CoT reasoning demonstrates strong performance in specific mathematical fields by leveraging step-by-step logical breakdowns to enhance problem-solving accuracy. Thus, while CoT provides a deeper, structured reasoning framework [Boye and Moell, 2025, Li et al., 2024a], it may struggle to satisfy the generalization requirements necessary for handling diverse mathematical challenges. The trade-off between GRPO’s structured logic refinement and CoT’s sequential deductive approach suggests that a hybrid reasoning mechanism may be essential for developing a more versatile and mathematically proficient LLM.
In summary, our findings align with prior research [Ahn et al., 2024a, Akella, 2024, Duan, 2025, Evstafev, 2025, Jiang et al., 2025, Li et al., 2024b, 2025, Mercer et al., 2025, Mirzadeh et al., 2024, Peng et al., 2024, Shao et al., 2024, Xin et al., 2024, Zhong et al., 2024, Zhou et al., 2024, Zhuang et al., 2024], further confirming that the recent invented model DeepSeek-R1 demonstrates strong mathematical reasoning compared to other LLMs. However, our research extends beyond previous work by providing a granular, field-specific analysis of LLM performance across different domains. We delve into the nuances of logical, arithmetic, and abstract reasoning capabilities, offering a holistic understanding of how different LLMs excel or struggle across mathematical disciplines. Despite its valuable contributions, our study has several limitations: 1). Limited Model Scope: We did not test other cutting-edge models, such as Qwen. 2). Restricted Benchmark Diversity. More importantly, future research should expand the scope of evaluated LLMs and incorporate additional benchmarks to gain a broader understanding of model strengths and weaknesses across varied mathematical disciplines. Notably, integrating GRPO with peer CoTs could allow models to leverage both structured logical optimization and stepwise deductive reasoning. This integration could lead to LLM swarms, where multiple reasoning frameworks synergize to enhance efficiency and accuracy. Meanwhile, our study reaffirms that distillation can significantly impair reasoning capabilities. Future efforts should explore adaptive distillation methods that preserve deep reasoning pathways while optimizing for computational efficiency.
5 Conclusion
Overall, our study presents a comprehensive comparative analysis of LLM performance in mathematical reasoning, identifying key strengths and weaknesses across multiple domains. We highlight the effectiveness of GRPO in structured problem-solving, demonstrating its ability to enhance logical consistency and rule-based reasoning. However, we also uncover its limitations in handling abstract mathematical concepts, where deep symbolic manipulation and theorem-based reasoning remain challenging for GRPO-trained models. Furthermore, our findings underscore the trade-offs associated with model distillation, revealing its potential drawbacks in preserving mathematical reasoning capabilities. Beyond performance evaluation, this study paves the way for future advancements in LLM-driven mathematical intelligence by outlining potential enhancements in reasoning frameworks, including hybrid models that integrate RL with structured step-by-step inference methods. By addressing these challenges, our work contributes to the development of next-generation AI systems capable of handling complex mathematical reasoning with greater accuracy, efficiency, and adaptability.
6 Acknowledgement
The authors gratefully acknowledge the support of High Performance Computing Services for providing the computational resources that contributed to the results presented in this publication/report.
Appendix A Appendix A
The table 2 summarizes the benchmark datasets used for evaluating the mathematical reasoning abilities of LLMs. The MATH dataset consists of free-form numeric problems drawn from math competitions, covering advanced topics like algebra, geometry, number theory, and combinatorics. GSM8K focuses on grade school arithmetic through structured word problems. The MMLU subsets include four college-level and high school-level categories—College Mathematics, Abstract Algebra, Formal Logic, and High School Mathematics—each formatted as multiple-choice questions and covering domains such as discrete math, algebraic structures, and logical reasoning.
MATH is derived from high-level mathematics competitions, including the AMC 10, AMC 12, and AIME [Hendrycks et al., 2021]. These problems are inherently more complex than standard K-12 mathematics tasks, requiring the application of non-trivial problem-solving strategies, logical inference, and domain-specific heuristics rather than direct formulaic computations. Meanwhile, MATH is widely regarded as a difficult benchmark, with LLMs achieving accuracy rates between 3.0% and 6.9% [Hendrycks et al., 2021]. Despite these low overall scores, LLMs demonstrate some degree of mathematical competency, as evidenced by: Up to 15% accuracy on the easiest difficulty level [Hendrycks et al., 2021]. The ability to generate step-by-step solutions that, while sometimes incorrect, remain coherent and contextually relevant. To gauge the dataset’s difficulty, human performance was also considered. For instance, a computer science Ph.D. student with no particular affinity for mathematics achieved approximately 40% accuracy. In addition, a three-time IMO (International Mathematical Olympiad) gold medalist attained 90% accuracy. Thus, these findings indicate that MATH is challenging for both human solvers and LLMs, making it an invaluable dataset for assessing and advancing mathematical reasoning capabilities in both LLMs and human problem solvers.
GSM8K, introduced in 2021, is a benchmark dataset consisting of 8,500 high-quality grade school-level math problems  [Cobbe et al., 2021]. Designed to incorporate high linguistic diversity while relying on fundamental mathematical concepts, it presents a unique challenge for state-of-the-art LLMs. Although the underlying math is relatively simple, the diverse problem formulations create significant hurdles, preventing many models from achieving consistently high accuracy. GSM8K is structured into 7,500 training problems and 1,320 testing problems, all carefully crafted by expert human problem writers. The problems primarily involve elementary arithmetic operations and typically require 2 to 8 logical steps to reach a solution. This dataset is widely used to evaluate logical reasoning and mathematical proficiency in LLMs and serves as a benchmark for various assessments, including the LLM Leaderboard. Notably, while some GSM8K problems are conceptually straightforward, they can still be challenging for even the most advanced LLMs, often exhibiting high variability in responses when the same problem is presented in slightly different ways. This highlights the ongoing difficulty of achieving robust mathematical reasoning in language models. In this study, we selected the 1.32K mathematical test set from GSM8K to systematically evaluate and compare the mathematical reasoning capabilities of each LLM.
The MMLU dataset is a comprehensive benchmark comprising multiple-choice questions from a diverse range of academic disciplines. Spanning 57 distinct tasks, it covers subjects in the humanities, social sciences, hard sciences, and mathematical reasoning, reflecting the breadth of knowledge that is essential for various fields of study [Hendrycks et al., 2020]. MMLU’s questions were manually curated by graduate and undergraduate students from publicly available sources, including practice questions from standardized exams such as the Graduate Record Examination (GRE) and the United States Medical Licensing Examination (USMLE). Additionally, it features questions designed for undergraduate courses and readers of Oxford University Press books. Notably, its mathematical reasoning component encompasses multiple subfields, categorized into areas such as "Abstract Algebra," "College Mathematics," "Formal Logic," and "High School Mathematics." For instance, questions in the "Abstract Algebra" category originate from professional mathematical practice, while the "High School Mathematics" category includes problems akin to those found in standard high school exams. This broad and structured dataset serves as a critical benchmark for evaluating the reasoning and problem-solving capabilities of LLMs.
@@ Appendix A Appendix A
Appendix B Appendix B
[…]